1. A certain planet has an average surface area that is 2.25% larger than the earth and a gravitational acceleration that is three fifth times than that of the earth’s gravitational acceleration. What is the velocity of escape on that planet? Ans. 8728.33 m/s
  2. Find the isogonal trajectories of the family of curves x2+y2=c if θ=45°. Ans. ln(x2+y2)+2 arctan (x/y) =k
1
Expert's answer
2021-11-25T15:15:32-0500

1.

For a spherically symmetric, massive body such as a star, or planet, the escape velocity for that body, at a given distance, is calculated by the formula


"v_e=\sqrt{\dfrac{2GM}{r}}"

where "G" is the universal gravitational constant, "M"  is the mass of the body to be escaped from, and "r" is the distance from the center of mass of the body to the object

An alternative expression for the escape velocity "v_e" is


"v_e=\sqrt{2gr}"

where "r" is the distance between the center of the body and the point at which escape velocity is being calculated and "g" is the gravitational acceleration at that distance.

Then


"v_e=\sqrt{2(\dfrac{3}{5})(9.81m/s^2)(\sqrt{1.0225}(6.4\times10^6m)}"

"\approx8728.33m/s"

2.

"x^2+y^2=c"

Differentiate both sides with respect to "x"


"2x+2y\dfrac{dy}{dx}=0"

"x+y\dfrac{dy}{dx}=0"



Replace "\dfrac{dy}{dx}" by "\dfrac{\dfrac{dy}{dx}-\tan 45\degree}{1+\dfrac{dy}{dx}\tan 45\degree}=\dfrac{\dfrac{dy}{dx}-1}{1+\dfrac{dy}{dx}(1)}" we get



"x+y\dfrac{\dfrac{dy}{dx}-1}{1+\dfrac{dy}{dx}}=0"

"x(1+\dfrac{dy}{dx})+y(\dfrac{dy}{dx}-1)=0"

"\dfrac{dy}{dx}=\dfrac{y-x}{y+x}"

Put "y=ux, \dfrac{dy}{dx}=x\dfrac{du}{dx}+u"


"x\dfrac{du}{dx}+u=\dfrac{ux-x}{ux+x}"

"x\dfrac{du}{dx}=\dfrac{u-1-u^2-u}{u+1}"

"\dfrac{u+1}{u^2+1}du=-\dfrac{dx}{x}"

Integrate


"\int\dfrac{u+1}{u^2+1}du=-\int\dfrac{dx}{x}"

"\dfrac{1}{2}\ln(u^2+1)+\tan^{-1}u=-\ln|x|+\dfrac{1}{2}k"

"\ln(x^2u^2+x^2)+2\tan^{-1}u=k"

"\ln(y^2+x^2)+2\tan^{-1}(y/x)=k"

which is the required trajectories of the given family of curves.